### v1

# Fast In-Spectrum Graph Watermarks

Abstract
We address the problem of watermarking graph objects, which consists in hiding information within them, to prove their origin. The two existing methods to watermark graphs use subgraph matching or graph isomorphism techniques, which are known to be intractable for large graphs.
To reduce the operational complexity, we propose FFG, a new graph watermarking scheme adapted from an image watermarking scheme, since graphs and images can be represented as matrices. We analyze and compare FFG, whose novelty lies in embedding the watermark in the Fourier transform of the adjacency matrix of a graph. Our technique enjoys a much lower complexity than that of related works (i.e. in $\mathcal{O}\left(N^{2}\log N\right)$), while performing better or at least as well as the two state-of-the-art methods.
I Introduction
Watermarking is the art of hiding information within a digital object. Watermarking has proven to be efficient in embedding robust information in complex data [18, 21]. Despite the explosion in the use of graph-structured data in various scientific fields [3, 7, 8, 5], the research of graph watermarking — especially for unweighted graphs — is still in its early stages [13, 38].
Creating graphs requires extensive data collection and pre-processing, making them highly valuable assets. Graphs emerge from a wide range of applications, including webs of trust in blockchains, circles in social networks, and co-purchasing networks in e-commerce platforms. The creators of these graphs often seek recognition, similar to the practices encouraged by Creative Commons licences. Watermarking schemes can satisfy this demand by embedding provenance information as a digital signature that can be extracted later to prove ownership.
Despite its potential, the watermarking of graph objects remains uncommon. None of the popular online platforms that share graph representations [24, 22, 32] currently propose watermarked graphs. This is mainly due to the limited number of existing techniques that are too computationally expensive. These techniques rely on intensive graph-related operations, such as subgraph matching and isomorphism [13, 38].
In this paper, we introduce a competitive watermarking scheme inspired by image watermarking techniques [4]. By treating adjacency matrices of graphs analogously to pixel matrices in images, we propose to leverage complexity-reduced computational operations for watermarking graph adjacency matrices. An overview of methods for watermarking graph-related objects is provided in the appendix B.
Contributions. Our contributions are as follows.
1) We introduce a novel analytical framework that applies an image processing scheme directly to the adjacency matrices of graphs, with the aim of using operations faster than the NP-complete related complexities of the related works [13, 38].
2) The main technical challenge in applying an image watermarking technique in a graph context is related to the binary nature of an edge’s presence in an adjacency matrix, as opposed to real numbers in an image matrix. This forces binarization. Our formal analysis studies the necessary binarization of the watermark to verify that it effectively leads to a unique watermark in the graph (i.e. without collision with another watermark insertion), setting the stage for the subsequent practical applications.
3) Finally, we perform a head-to-head comparison of FFG with the related work in the graph domain (schemes by Zhao et al. [38] and Eppstein et al. [13]). We conclude that FFG performs at least as well as its competitors, with the advantage of being much faster.
II Goals and Threat Model
The following section describes watermarking schemes and the goals of such methods.
II-A Watermarking Graphs: Threat Model
A graph watermarking scheme is a set of functions (Keygen, Embed, Extract), applied to a graph [13]. With Keygen, the owner of a graph generates a secret key. This key is embedded into her graph $G$ with the Embed operation. It produces a modified graph $G_{W}$ and a watermark $W$ (which is related to the difference between the original and the watermarked graph). The graph’s owner can prove her ownership of a suspected shared graph $G^{*}$ with the watermark $W$ using the Extract function. Extract returns a success if $G^{*}$ is watermarked with $W$. Otherwise, either 1) $G^{*}$ was not watermarked with $W$ and the considered watermarking scheme (or not watermarked at all), or 2) an attacker who wanted to prevent extraction has modified $G^{*}$ enough to prevent Extract from succeeding.
II-B Design Goals for Watermarking Schemes
The goals of a watermarking scheme are as follows [38]:
(Goal 1) Low distortion
The watermark embedding with Embed must have a low impact on the original graph to preserve the intrinsic value of that graph.
(Goal 2) Watermark uniqueness
The Embed function must be injective. In other words, a graph watermarked independently with two different keys should not produce two identical graphs.
(Goal 3) False positives and negatives
A good watermarking scheme must minimize false positives without generating too many false negatives.
(Goal 4) Robustness to modifications
A robust watermarking scheme must be resistant to modifications (i.e. attacks) performed on the watermarked graph. More specifically, a key extraction (Extract) must work with a high probability, despite attacks.
Other goals such as undetectability and distortion effect are defined and studied in Appendix E. These are additional goals we set with respect to the state-of-the-art watermarking methods.
III FFG: Cox et al. Scheme for Graphs
This work considers undirected and unweighted graphs, in order to compete with state-of-the-art schemes [13, 38] described in Section V-A. Since the goal is to exploit watermarking schemes originally designed for images, we assume that we can label the input graph arbitrarily, so that a deterministic mapping of its vertices in the adjacency matrix can be performed [14, 17]. This assumption is realistic in many cases where vertices are naturally labeled and it is consistent with the fact that for images, pixels have a deterministic position in their matrix representation.
III-A Framework Rationale
Notations introduced hereafter are also summed up in Table III and Figure 7 (Appendix A).
[…]
Therefore, our framework includes additional steps to transform $A^{\prime}$ back into an undirected and unweighted graph. Part of our challenge is to analyze the consequences of these steps. More specifically, the required binarization is the thresholding of all values in $A^{\prime}$ by the average of $A$. This real value is denoted $av(A)$. Each element is set to 00 if its modulus is less than the average of the original adjacency matrix $A$, otherwise $1$.
Our framework yields a proper graph representation through its new adjacency matrix $A_{W}$. From this matrix, the resulting graph can be shared at will by the owner of the graph.
Next, we describe the original scheme of Cox et al. for watermarking images, before detailing its adaptation to graphs in our framework.
III-B Watermarking Images with the Cox et al. Scheme
In [4], the watermark is a sequence of a fixed number of random reals chosen from a Gaussian distribution. The authors compute the Fourier transform of the original image and insert the watermark into the largest low-frequency coefficients of the transformed matrix. They propose three ways to insert the watermark into the highest magnitude coefficients, including a simple additive method.
For a given image (possibly modified), a watermark is extracted in the spectral domain and compared with the original watermark. A similarity score is computed, which decides whether this image is watermarked or not.
III-C The FFG Scheme for Watermarking Graphs
We now propose FFG as an adaptation of the Cox et al. scheme to be applicable to unweighted graphs. We first present the three functions Keygen, Embed and Extract. The pseudocode is given in Algorithm 1. We then discuss the main differences with the original Cox et al. scheme.
Algorithm 1 FFG scheme
1:function Keygen($m,\sigma$)
2:    $\omega=\left[\mathcal{G}(0,\sigma^{2}),\ldots,\mathcal{G}(0,\sigma^{2})\right]$triangleright of length $m$
3:    Return $\omega$
4:function Embed($A,\omega$)
5:    $m=|\omega|$
6:    $th=av(A)$triangleright real number
7:    FT(A) = Fourier transform of $A$
8:    $\chi=\texttt{argsort}(|FT(A)|)[0:m]$triangleright argsort returns the indices that would sort its argument
9:    $\mathcal{W}[i,j]=\left.\begin{matrix}\omega[n]\text{ if there exists }n\leq m,\;\chi[n]=(i,j)\\ \end{matrix}\right.$
10:    $A^{\prime}=A+FT^{-1}(\mathcal{W})$
11:    $A_{W}[i,j]=\left\{\begin{matrix}1\text{ if }|A^{\prime}[i,j]|>th\\ \text{else }0\end{matrix}\right.$triangleright binarization
12:    Return $A_{W}$
13:function Extract($A$, $A^{*},\omega,\theta$)
14:    $A_{W}=...$triangleright generated as in Embed
15:    $W^{*}=FT(A-A^{*})$
16:    $W=FT(A-A_{W})$
17:    $s=\left\|W^{*}-W\right\|_{2}$
18:    Return $s\leq\theta*\left\|W\right\|_{2}$triangleright True: watermark retrieved
Key generation
Keygen$(m,\sigma)$ generates a key as a Gaussian vector of $m$ values with standard deviation $\sigma$. This step is identical to that of Cox et al..
Watermark embedding
The graph owner wants to embed a key $\omega$ generated with the function Keygen into her graph $G$ thanks to its adjacency matrix $A$. First, the sequence $\chi=\{(i_{1},j_{1}),\dots,(i_{m},j_{m})\}$ is computed as the sequence of the first $m$ indices that would sort the modulus of the Fourier coefficients of $A$. The sequence $\chi$ is used to compute the intermediate watermark $\mathcal{W}$ of the graph owner in the Fourier spectrum. $\mathcal{W}$ is the matrix representation of the key where each element $\omega[n]$ is at position $\chi[n]=(i_{n},j_{n})$. Then, the Fourier inverse of the intermediate watermark $\mathcal{W}$ is added to the adjacency matrix $A$ to obtain a real matrix $A^{\prime}$. The algorithm continues with binarization and symmetrization to obtain $A_{W}$. The watermark $W$ is defined as the difference of all coefficients of the Fourier transform of $A$ and $A_{W}$: $W=FT(A)-FT(A_{W}).$
Watermark extraction
The suspected watermark $W^{*}$ is compared to the original watermark $W$ using the $2$-norm of their difference. If the difference between $W^{*}$ and $W$ is less than a threshold, the extraction is hypothetically successful.
The specificity of FFG
The first difference between the original scheme of Cox et al. [4] and FFG is that the watermark key cannot be chosen arbitrarily, but instead depends on the binarization operation. What is left after binarization depends on the chosen length $m$ and the standard deviation $\sigma$ of the elements in that key. Recommendations in Cox et al. [4] are no longer valid due to this binarization step. Section IV-C discusses how to properly obtain a key.
The second specificity is also due to the binarization step. This step causes the key to be spread over the entire transformed matrix. As a consequence: $\mathcal{W}\neq W.$ Thus, the watermark is defined over all matrix coefficients and not only over the ones specified by $\chi$. The function Extract is applied to the watermark $W$ resulting from the binarization after embedding $\omega$ with the Cox et al. scheme, and not applied directly to the key $\omega$ in the coefficients specified by $\chi$. Given the key $\omega$, the watermark $W$ can be derived by the embedding function or by calling Extract before comparing to $W^{*}$.
III-D Scheme complexity vs related works
[…]
For comparison, the scheme from Zhao et al. leverages the NP-complete subgraph matching routine. They argue that the computational complexity is lower on graphs with a ”very high level of node heterogeneity” in terms of degrees, without providing a complexity analysis for this claim (see [38], Section 4); the problem still remains NP-complete in the general case. While Eppstein et al. [13] prove the correctness of their scheme for Erdős-Rényi and power-law graph families, they do not provide a bounded complexity analysis for these, and thus also remain in the NP-complete complexity class due to the subgraph isomorphism routine. All three schemes are compared in terms of their practical execution times in Section V-B.
IV Experiments: Compliance to Design Goals
This section presents the results of an extensive experimental study of FFG. It is organized to shed light on the practical compliance of FFG to the design goals detailed in Section II-B.
IV-A Experimental Setup
Experiments are performed using Python 3. Reported timings are obtained from execution on a server equipped with Intel Xeon E5-2630 v3 CPUs, running at 2.40GHz.
Graph generative models
For experiments that require the variation of a graph property, graph generators from the Python package NetworkX [16] are used. These are well-known graph models, that have been used to model various real-world graphs: Erdős-Rényi (hereafter referred to as ER), Barabási-Albert (BA), or Watts-Stroggats (WS) models.
Large and real graphs
Experiments with FFG are also performed on eight real-world large graphs with different edge densities, obtained from the SNAP repository [24] (e.g. Pokec) and the NetworkRepository [32] (e.g. Flickr). These graphs have different structures, which makes them interesting to study.
Evaluation metrics
The properties and information to preserve are highly context-dependent, depending on the intended use of a graph. It is impractical to cover all possible scenarios. Due to its generality and widespread use in the state-of-the-art [13, 38], we consider the graph edit distance, as a metric to quantify the degradation of a graph after these attacks:
Definition.
The graph edit distance (denoted ED) is the percentage of edges that distinguish two graphs. In the sequel, the distance is measured relative to the original graph.
For instance, if the $embed$ function flips $200$ edges from a graph $G$ containing $2,000$ edges, then the distance based on the fraction of edges between $G^{\prime}$ and $G$ is $10\%$ edges. We emphasize that the ED can be greater than $100$% because we can add edges and not just remove them. In fact, one can flip at most $N(N-1)/2$ edges from the input graph.
IV-B Adaptations from the Theoretical Scheme
IV-B1 FFG Scalability
To increase the scalability of their schemes when applied to large graphs, both related works [13, 38] filter out the input graph to operate on a smaller set of vertices. Inspired by the Cox et al. scheme [4] FFG also performs a dimensionality reduction. This step is described in detail in the Appendix D.
We empirically evaluate the scalability of FFG by increasing the size of the input graph from $N=10^{3}$ to $10^{7}$. To maintain a similar graph structure as the size increases, a BA graph generator is used. Its attachment parameter is set to $3$ (i.e. each new vertex attaches to $3$ existing vertices according to preferential attachment). The length of the key is arbitrarily always the same ($m=200$) because all the graphs have the same attachment parameter (see Section IV-C).
The measured timings are as follows. For embedding: $0.20,~{}18,~{}20,$ $46$ and $358$ seconds for BA with sizes $1k,~{}10k,~{}100k,~{}1M$ and $10M$ respectively. For extraction, and the same graph sizes: $0.13,~{}15,~{}16,~{}17$ and $20$ seconds. The execution time of the key generation is almost constant at $0.8$ milliseconds. These observations are fully consistent with the timing complexity analysis, and therefore confirm that our FFG is tractable for large graphs.
IV-B2 Resilience to Attacks and the Similarity Threshold
As in the original Cox et al. scheme and in the related work [38] and [13] for graphs, there is a threshold parameter $\theta$ driving the success of the extraction function in Algorithm 1. The following paragraphs deal with this parameter $\theta$.
The ideal case: the watermarked graph is never attacked.
We demonstrate that if attacks on $G_{W}$ are not considered, $\theta=0$ can be set with the guarantee that the watermark of the graph’s owner will always be extracted. That is, considering unattacked watermarked graphs, no false negatives can occur.
Proof. A false negative occurs if Extract returns a failure even though the tested graph was indeed watermarked with a given watermark. In particular, the extraction will fail only if the following condition is not satisfied: $\left\|W^{*}-W\right\|_{2}\leq\theta*\left\|W\right\|_{2},$ where $W^{*}$ and $W$ are computed according to Algorithm 1, and while the tested graph $G^{*}$ is the watermarked graph $G_{W}$, that has not been attacked. Thanks to the deterministic mapping of vertices to adjacency matrices, $A^{*}=A_{W}$. Since the Fourier transform is linear, the condition $\left\|W^{*}-W\right\|_{2}\leq\theta*\left\|W\right\|_{2},$ becomes: $0\leq\theta*\left\|W\right\|_{2}$, which always holds since $\theta\geq 0$. Thus if $\theta=0$, no false negatives can occur with the FFG scheme.
A realistic case where the watermarked graph is attacked
In a real scenario, after the owner of a graph shares her watermarked graph, it can be modified. A false negative occurs when Extract returns a failure, despite the tested graph was indeed watermarked with the owner’s watermark. In this case, $\theta>0$ defines the similarity above which the graph’s owner considers the extracted watermark to be close enough to the original one, despite modifications.
We set $\theta$ to tolerate $10$% edge flips, which is clearly a conservatively high value (according to related work [13, 38], as and it will be illustrated in Section IV-C). Consequently, no false negatives can occur with less than $10$% edge flips. The effect of $\theta$ on false positives is studied in Section IV-E.
IV-C (Goal 1) Low Distortion
We now experiment with the distortion caused by embedding the watermark in a given graph. We study it theoretically in Appendix C.
Practical guidelines for low distortion
From Section II-B, low distortion is targeted for the watermark embedding. According to related work [13, 38], the ED of large graphs must be at most $10^{-2}$edges. We now give a method to automatically set the parameters suitable for watermarking with such a minimal distortion.
Figure 1 illustrates the link between $\sigma$ and $m$ for our three generative graph models, and a small distortion of $0.005\%$ edges.
Key lengths used in [13] depend on densities (being defined as the number of edges over the number of vertices) [26], while those of [38] depend on the number of vertices. We set the key length $m$ of FFG as the key length used in [13] because the experiment in Section IV-D shows that FFG is also related to densities.
Once the key length $m$ is set as in [13], it is always possible to find a value of $\sigma$ to watermark the graph while achieving the distortion goal. Experiments have shown that for any graph $G$ and for any fixed key length $m\in[\![1,N_{0}*(N_{0}-1)/2]\!]$, there exists a value $\sigma_{max}$ to watermark the graph with a strictly positive edit distance. An automated way to set $\sigma$ under the small ED constraint is to use a dichotomous search in $[\![1,\sigma_{max}]\!]$. As shown in Appendix D, the proposed guidelines do not affect the scalability of FFG.
Figure 1: The relation between $m$ and $\sigma$ to watermark graphs with an ED ]0,0.005[$\in$]0,0.005[∈ ] 0 , 0.005 [ % of edges, for the three graph models.
Graph $m$ $\sigma$ Watermarking ED
BA $210$ $1,750$ $3.10^{-8}$
Flickr $3,250$ $32,000$ $2.10^{-8}$
Pokec $170$ $7,000$ $5.10^{-9}$
TABLE I: Parameters and resulting distortion (ED).
The three large and real graphs are watermarked using this method, leading to the parameters in Table I. Note that the values of $m$ differ significantly between graphs, depending on their inner topological properties. The resulting EDs are effectively kept well below the target of $10^{-2}$%. Goal 1 is thus achieved.
Further experiments on the impact of FFG on graphs are studied in Appendix E.
IV-D (Goal 2) Uniqueness of Watermarks
To assess the uniqueness property resulting from FFG, two keys $\omega_{1}$ and $\omega_{2}$ are generated with the same parameters ($m$ and $\sigma$) found by the automatic dichotomous procedure. There is a lack of uniqueness if the two keys result in the same watermark (i.e. if $\omega_{1}\neq\omega_{2}$ and $G_{W_{1}}=G_{W_{2}}$).
We experiment with uniqueness for the three graph models and various graph densities in Figure 2. There are some collisions for the lowest densities, beyond $2$. This is not an issue per se, as these graphs are rare or atypical: a density less than $2$ means that a graph has, or is close to having, disconnected components. Above this threshold, there is no collision; this is the expected setup, since collected graphs generally have high densities (e.g., from 12 to 36 in web crawls [26], or without density limit for graphs studied in node classification, where edges are even removed to reduce the density before processing [25]). Goal 2 is thus validated for practical setups.
Figure 2: Success rate ($y$-axis) in inserting a unique watermark in three graph models, for 1M vertices graph models and different densities on the $x$-axis.
IV-E (Goal 3) Low False Positives and Negatives
We now study the effect of $\theta$ on false positives.
Let $G_{W}$ be the watermarked graph of $G$ with the key $\omega$, and $G^{*}$ be any other graph not watermarked with $\omega$. Recall that Algorithm 1 with $G^{*}$ results in a false positive if the extraction of $W$ in $G^{*}$ succeeds.
Figure 3: The impact of $\theta$ on extraction false positives (BA graphs).
The goal of the following experiment is to determine the value of $\theta$ for which false positives occur when attempting an extraction on non-watermarked graphs. Ten BA graphs are generated with density $\frac{|E|}{|V|}=5$ (according to the uniqueness experiment in Figure 2). For each of them, we first inserted a key $\omega$; then we re-generated five more graphs with the same generative parameters, and tried to extract $\omega$ from them. Extractions are performed for different values of $\theta$. Figure 3 shows that as long as $\theta$ is less than or equal to $5$, there are no false positives. Then, values of $\theta$ above $6$ introduce some. Thus, $\theta=1.4$ is consistent with the $2$-norm of the watermark to tolerate $10$% edge flips for a graph of density $5$ in BA models. Goal 3 is then achievable by parameterization.
In the remaining experiments, $\theta$ is set to tolerate $10$% edge flips since no false positives or false negatives occur in this setup.
IV-F (Goal 4) Robustness to Attacks
IV-F1 Intensity and Impact of the Attack
In order to assess a plausible intensity of an attack that a graph can sustain without losing its utility, we show in Figure 4 a simple function performed on the three first graphs of Table II: the ranking of the top-$1000$ nodes according to the degree centrality [6]. The figure plots the changes in this ranking of the most central nodes, according to the classic Spearman metric [10]. It measures the correlation between two rankings of the top nodes (on the original graph and on the attacked version). The correlation drops dramatically as the attack increases, with even the low value of $1\%$ of edges flipped. Thus, it appears that this random attack disrupts the connectivity of the central nodes, clearly questioning the usefulness of the resulting graph. In this context, $10\%$ of flipped edges is an extreme attack that appears to remove the utility of the graph, and can thus be considered a very conservative upper bound.
[…]
IV-F2 Robustness
The performance of FFG is reported on real large graphs in Table II.
The key lengths are similar to those used in Eppstein et al. [13] and standard deviations are set to target watermarks that cause little distortion. Despite their various densities, all of them were watermarked with indeed small distortions (EDs of $10^{-6}$ at the very least, most often at $10^{-8}$, and even as low as $10^{-10}$ for the graph ”inf-belgium-osm”).
Graphs $|V|$ $\frac{|E|}{|V|}$ $m$ $\sigma$ ED $\theta$ $Suitability$
inf-belgium-osm 1.4M 1.1 54 250 $3.10^{-10}$ 0.6
soc-YouTube-ASU 1.1M 2.6 200 12,000 $1.1.10^{-8}$ 2
hollywood-2009 1.1M 52.7 162 224,000 $1.6.10^{-8}$ 205
rgg-n-2-20-s0 1M 6.6 119 1,624 $3.8.10^{-8}$ 10.2
kron-g500-logn20 1M 42.6 71128 16,000 $3.5.10^{-8}$ 374
scale21-ef16-adj 1.2 M 51 44,743 12,000 $1.7.10^{-8}$ 233
roadNet-PA 1.1M 1.4 48 184 $6.10^{-6}$ 0.6
delaunay-n20 1M 3.0 63 61 $4.3.10^{-8}$ 6.2
TABLE II: Experiment settings and results for FFG, on real large graphs.
We now subject these graphs to attacks and observe the resilience of FFG. The last column $Suitability$ shows that the extractions are successful despite the attack. We also note that additional tests have shown that all watermarks can as well be extracted even under an attack that flips $100$% edges, for a value of $\theta$ set to be resistant to only $10$ edge flips. This underscores the robustness of FFG, and the reach of Goal 4.
This experimental section confirms both the applicability of FFG to real graphs, the low distortion of the scheme on these graph structures, and its resilience to significant attack strengths.
IV-F3 Densities and the Resulting Threshold
The last experiment for FFG is to test its robustness to different graph densities, for the three graph models. We take the three graph generators (BA, WS, ER) and vary the density $\frac{|E|}{N}$ of graphs to $5$, $30$ and $50$. $5$ is a low density which respects the uniqueness condition, $30$ and $50$ are medium and high densities according to [26] and the graphs taken from SNAP [24]. For each generator and each density, we set $\theta$ to be resistant to $x$edge flips ($x$-axis). This resilience is observed on $3$ runs per point on Figure 5, where $\theta$ is set by a dichotomous search for the effective resilience to the attack.
Figure 5: The choice of $\theta$ in FFG to be resistant to $x$edge flips depending on the density and the type of graph model.
Figure 5 shows that the resulting threshold $\theta$ depends primarily on the type of graph. For BA or WS graphs, a setting $\theta=1$ is sufficient for FFG to be resistant to attacks with up to $10\%$ edge flips ($x$-axis), regardless of the density of the original graph. Density is the most relevant parameter when watermarking ER graphs, where $\theta$ can vary from $0.3$ to $5$ to be resistant to attacks of up to 10% edge flips.
V A Comparative Benchmark
Finally, we perform a head-to-head comparative benchmark of FFG with the two state-of-the-art approaches [38] (Zhao) and [13] (Eppstein), both in terms of running time and resilience to attacks.
V-A Related work
The scheme by Zhao et al. [38]
There are four steps in the graph embedding function of Zhao et al. [38]. First, they generate a random seed based on cryptographic keys. Then, they generate a random watermark graph (Erdős-Rényi) using the seed and match it with a subgraph of $G$. The subgraph selection is based on the structure of the vertices (i.e. their neighborhoods). Finally, the embedding is performed by an XOR operation between the subgraph and the random watermark graph. For the extraction, they regenerate the watermark graph and the subgraph like in the embedding process. After that, they identify in the potential $G^{*}$ which vertices are candidates to match the subgraph (again using the vertex structure). Finally, they extract the watermark using a recursive algorithm. Since the code was not provided by the authors, we implemented it in Python 3. We note that our execution results are consistent with those presented by the authors in their experimental section (Figure 2b in [38]). We set $\theta=0.95$ and $B=25$, for an acceptable execution time.
The scheme by Eppstein et al.  [13]
The scheme of Eppstein et al.[13] consists of the same three functions Keygen, Embed and Extract as the FFG scheme. In the scheme of Eppstein et al., private keys generated by the Keygen function are sets of vertex pairs in a graph containing as many vertices as in the graph to be watermarked. There are three main steps in the embedding function. The first step is ordering and labeling the vertices of the graph, considering degrees and using bit vectors. The vertices are divided into three categories: high, medium, and low degree vertices. Then, each edge contained in the private key is randomly flipped according to the probability of its statistical existence in the graph. The function Extract uses graph isomorphism to identify the watermarked graph. Unlike FFG and Zhao [38], this method identifies a graph between multiple (simultaneously created) watermark copies, not just a single watermark graph with a given key. The selected copy is the closest copy considering the Hamming distance.
V-B Execution Timings
We first benchmark the runtime of the Embed and Extract functions of the competing watermarking schemes. The keys are generated according to the recommendations in each paper; their timings are not displayed, as they are all very fast. Scalability improvements suggested by each paper have been implemented.
(a) Execution timings
(b) Robustness
Figure 6: Benchmark comparison between FFG, Eppstein [13], and Zhao [38]. (a) Timing ($y$-axis) of the Embed and Extract functions for the three competing schemes, depending on the BA graph size $N$ ($x$-axis). Executions taking less than the timeout (fixed at 25 minutes) are kept and plotted. (b) Robustness of the three schemes ($y$-axis) facing attacks: the normalized number of successful extractions, facing increasing edge flips (measured by edit distance on the $x$-axis as ratio of edges) on three large graphs.
In Figure 6a, timings are measured in seconds as a function of the number of vertices $N$ in an increasing size BA graph (from $1k$ to $10M$ vertices, with an attachment parameter of $3$). We set a timeout of $25$ minutes for the experiments (dotted gray horizontal line), so that all scheme executions –each running on a single core– are discarded if they exceed this threshold. For small and medium-sized graphs ($N\leq 10^{5}$), the embedding function of Zhao et al. is the most efficient method, while its extraction takes the longest. For medium and large graphs (more than $10^{5}$ vertices), the FFG scheme outperforms the state-of-the-art. Without exceeding the timeout, FFG can watermark graphs with $10$ times more vertices than the scheme of Eppstein et al. and up to $100$ times more vertices than the scheme of Zhao et al.
Note that extractions are performed on a non-attacked $G_{W}$, which is the most favorable scenario for the schemes of Eppstein et al. and Zhao et al.: their extractions are longer when applied to attacked graphs, because the approximate subgraph matching is harder (while the matrix representation in FFG makes it time-invariant to attacks). The extractions of Zhao et al. and Eppstein et al. are linear in the number of vertices, while FFG is almost constant after $10k$ vertices.
Even though Zhao et al. and Eppstein et al. also perform dimensionality reduction, they use approximate subgraph isomorphism in the extraction, which is computationally expensive, if not intractable (NP-complete) for increasingly large graphs. FFG relies on vertex labels (in parallel with images) to avoid such a tedious operation and then works significantly faster.
V-C Robustness to the Edge Flip Attack
In Table II, we observed that FFG is resistant to $100$% edge flips on real graphs. (Recall that $100\%$ is relative to the number of edges in the original graph, which here means that $|E|$ edges were flipped among the $N(N-1)/2$ possible flips.) Since Zhao and Eppstein schemes can at best achieve this resilience, we now test the watermark extractions of the three schemes, despite attacks on three watermarked graphs.
Figure 6b shows the results for the three large graphs. The success rate is shown on the $y$-axis. The $x$-axis shows the number of edges flipped, measured by the ED. Each experience is run three times (the standard deviation is zero, as there is no variation in the results across all the performed executions).
We observe that FFG performs better than the scheme of Zhao et al. for the three graphs. The scheme of Zhao et al. fails before $10$% edges are flipped, while FFG handles flips equivalent to $100\%$ edges. FFG also beats the Eppstein scheme, except on the Flickr graph where Eppstein survives $1000\%$ of flips.
We conclude that FFG is at least as robust –except on one configuration–, if not better, against random edge flips than its two competitors. Overall, the high robustness of FFG (resilience $\geq 100\%$ of flipped edges, see Table II and Figure 6b) makes it practical even in scenarios with extreme attack strengths.
VI Conclusion
Large and real graphs are valuable assets that can be tracked for provenance using watermarking techniques. This paper bridges the gap between the image processing and the graph domains. While previous state-of-the-art schemes relied on standard graph manipulations, we have shown that knowledge and techniques from the multimedia community can be ported to the graph domain, yielding significant complexity improvements. This is achieved by studying the transitions between a graph and its real-valued image-like representation, and vice versa. Our method FFG shows strong performance on several metrics, including low false positives and negatives.
For future work, our framework can be extended with other binarization methods, such as the optimized thresholding method proposed by Otsu [27]. This may lead to even smaller watermarks, which is crucial for minimizing distortions in the watermarked graph. Another promising direction is to adapt recent image watermarking schemes to our framework.
Appendix A Notations
Table III is a summary of the main notations used in our paper. In particular, the differences between $\omega$, $\mathcal{W}$ and $W$ are noted. It is complemented by Figure 7, which illustrates our method.
$G$ An arbitrary undirected and unweighted graph
with $N$ vertices $V$, and edge set $E$
$G_{W}$ The watermarked graph
$G^{*}$ Another arbitrary graph to test for the watermark presence
$A$, $A_{W}$ Adjacency matrices of graph $G$ and $G_{W}$ respectively
$A^{\prime}$ Matrix obtained before the binarization of $A_{W}$
$\theta$ Similarity threshold for the watermark extraction success
$\omega$ Watermark key, of length $m$
$\sigma$ Standard deviation for generating the key
$\chi$ List of indices to insert the key
$\mathcal{W}$ Intermediary (not yet binarized) watermark
$W$ Final watermark
TABLE III: Main notations in this paper.
Figure 7: The FFG framework for watermarking unweighted graphs using an image domain scheme (hereby Cox et al. [4]).
Appendix B Watermarking non-graph objects
This section is an overview of methods for watermarking graph-related objects that are not graphs themselves.
Watermarking various digital objects
Initially introduced to protect multimedia objects [35, 4], watermarking techniques are now being applied to a wide range of other areas, such as the model protection of deep neural networks [23], graph neural networks [37] or large language models [20] to just cite a few recent works. Watermarking approaches sometimes use graph abstractions to hide information in the object, as it is the case with software watermarking. (See e.g. [9] for a review.)
In software watermarking approaches, graph abstractions capture the interactions between variables of a source code to be watermarked. The purpose is thus to watermark the software; graph representations are simply leveraged to build watermarking schemes, and are not the target of the watermarks. Some approaches focus on databases, to watermark datasheets or RDBMS tuples, for example [1, 5, 21]; directly watermarking graph objects remains much more general, as targeting graph structures themselves [38, 13], and because no assumption is made on the presence or absence of tuples and rich data to in which embed information into.
In graph signal watermarking [2], since signals are data defined on the vertices of graphs, it is natural to watermark them by using graph eigenvectors. However, this technique thus consists of watermarking signals defined over graphs, and it never watermarks graphs themselves.
Watermarking graphs
This paper is dedicated to watermark unweighted graphs. Recently, two graph watermarking approaches have been introduced that use purely graph techniques [13, 38] to watermark graphs. Since they are direct competitors to FFG, both schemes are described in detail in the main part of the paper (Section V), and then benchmarked.
Digital image watermarking
Efficient image watermarking has motivated decades of research. Images can be watermarked by taking into account features such as the perception of the human visual system (HVS) [34], or by applying structural operations to them such as JPEG compression [36]. Ingemar J. Cox et al. [4] proposed the general method of Fourier transform and in-spectrum watermarking. This work is still used today for other types of transforms such as wavelets [12] or fractional Fourier transforms [11].
Appendix C Theoretical Analysis of Uniqueness
This section aims to complete the experimental study of FFG. In particular, with respect to the uniqueness and success of a watermark.
We have defined watermarking schemes and how Cox’s watermarking scheme can be adapted to graphs. We now study the effect of binarization, which is related to design goal 2.
C-A Uniqueness of Watermarks
Let $G_{W_{1}}$ be the watermarked graph resulting from a pair $(G,\omega_{1})$. $G$ is the graph to be watermarked and $W_{1}$ the resulting watermark from $\omega_{1}$. The uniqueness problem is to determine whether $G_{W_{1}}$ is unique.
The only source of randomness in the scheme is the generation of the key; thus if $G$ is watermarked with another key $\omega_{2}$ resulting in $G_{W_{2}}$, then there is uniqueness if $G_{W_{1}}\neq G_{W_{2}}$. This property is not trivial because of the binarization which spreads keys over all the Fourier coefficients of the adjacency matrices.
We then study the probability of uniqueness:
$P(G_{W_{1}}\neq G_{W_{2}})=1-P(G_{W_{1}}=G_{W_{2}})$ (1)
Proof
Let $A_{W_{1}}$, respectively $A_{W_{2}}$, be the adjacency matrices of $G_{W_{1}}$, respectively $G_{W_{2}}$. By the hypothesis of the deterministic mapping of vertices into adjacency matrices: $P(G_{W_{1}}=G_{W_{2}})=P(A_{W_{1}}=A_{W_{2}}).$ Similarly, let $A^{\prime}_{1}$, respectively $A^{\prime}_{2}$ be the matrices corresponding to $A_{W_{1}}$, respectively $A_{W_{2}}$ before the binarization.
By definition, binarization is done using the average of the adjacency matrix $A$ of $G$:
$P(G_{W_{1}}=G_{W_{2}})=P\left(\bigcap_{i,j}(A_{W_{1}}[i,j]=A_{W_{2}}[i,j])\right)$
$=P\left(\bigcap_{i,j}\begin{matrix}[A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=A[i,j]=0\\ \cup A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=A[i,j]=1\\ \cup A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=1-A[i,j]=0\\ \cup A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=1-A[i,j]=1]\end{matrix}\right)$
$=P\left(\bigcap_{i,j}\begin{matrix}|A^{\prime}_{1,2}[i,j]|\leq av(A)\wedge A[i,j]=0\\ \cup|A^{\prime}_{1,2}[i,j]|>av(A)\wedge A[i,j]=1\\ \cup|A^{\prime}_{1,2}[i,j]|\leq av(A)\wedge A[i,j]=1\\ \cup|A^{\prime}_{1,2}[i,j]|>av(A)\wedge A[i,j]=0\end{matrix}\right).$ (2)
Here the notation $A^{\prime}_{1,2}$ means that the result must be true for both $A^{\prime}_{1}$ and $A^{\prime}_{2}$ respectively and $av(A)$ is the average of all elements of the matrix $A$.
By the linearity of the inverse Fourier transform,
for $n\in\{1,2\}$:
$A^{\prime}_{n}=A+FT^{-1}(\mathcal{W}_{n}).$ (3)
with
$FT(A)[i,j]=\sum_{k=1}^{N}\sum_{l=1}^{N}A[k,l]e^{-2\imath\pi\left(\frac{ik}{N}+\frac{jl}{N}\right)}$
and
$FT^{-1}(A)[i,j]=\frac{1}{N^{2}}\sum_{k=1}^{N}\sum_{l=1}^{N}A[k,l]e^{2\imath\pi\left(\frac{ik}{N}+\frac{jl}{N}\right)}.$
By construction, all the elements of each key $\omega$ (and its matrix version $\mathcal{W}$) are independent. The probability $P(G_{W_{1}}=G_{W_{2}})$ in Equation (1) can be computed with the cumulative distribution of $FT^{-1}(\mathcal{W}_{n})$ ($n\in\{1,2\}$) in Equation (3).
C-B Computing the Cumulative Distribution
First, note that:
$FT^{-1}(\mathcal{W}_{n})=Z\mathcal{W}_{n}Z,$
where $Z[k,l]=\frac{1}{N}e^{2\imath\pi kl/N}$. This shows that $FT^{-1}(\mathcal{W}_{n})$ is a matrix version of a Gaussian vector. With this, we can get an exact formula for the probability distribution function of this inverse Fourier transform flattened to one dimension [15]. In fact, let $\tilde{\mathcal{W}_{n}}$ be the flattened vector of $\mathcal{W}_{n}$, where:
$\forall k\in\{1,...,N^{2}\},\tilde{\mathcal{W}_{n}}[k]=\mathcal{W}_{n}[\lceil k/N\rceil,(k-1)\mod N+1],$
where $\lceil k/N\rceil$ being the upper integer part of $k/N$ and $(k-1)\mod N$ being the rest of the Euclidean division of $k-1$ by $N$. Using these notations, we easily show that:
$FT^{-1}(\tilde{\mathcal{W}_{n}})=\tilde{Z}\mathcal{W}_{n},$
where $\tilde{Z}=Z\bigotimes Z$ is the $N^{2}\times N^{2}$ Kronecker product of $Z$ with itself. Let $(z_{k,l})$ be the coefficients of $Z$, then for $(k^{\prime},l^{\prime})\in\{1,...,N^{2}\}$:
$(\tilde{Z})_{k^{\prime},l^{\prime}}=z_{\lceil k^{\prime}/N\rceil,\lceil l^{\prime}/N\rceil}z_{(k^{\prime}-1)\mod N+1,(l^{\prime}-1)\mod N+1}.$
Since $\tilde{\mathcal{W}_{n}}$ is a Gaussian vector, $FT^{-1}(\tilde{\mathcal{W}_{n}})$ is also a Gaussian vector. Its average and standard deviation are computed using the matrix $\tilde{Z}$. Therefore, we can use standard approximation techniques to approximate Equation (C-A).
This section has successfully derived a precise analytical form for evaluating the probability of a watermark’s uniqueness. However, considering the size of the vectors involved, large graphs will require a significant approximation time. In practice, direct experiments such as those in section IV-D will be preferred. Despite this computational challenge, the analytical insights shed light on the theoretical components at stake in discretizing a watermark in an unweighted graph.
Application on large graphs.
The connection between graph densities and collisions is given by the theoretical analysis in Equation (C-A).
Indeed, as $av(A)=\frac{|E|}{N^{2}}$, the threshold for binarization is equal to the density divided by $N$. For the $N=1M$ graphs (density less than $10$) in Figure 2 and the given keys:
$av(A)\simeq 0$
$avg(|FT^{-1}(W)|)\simeq 0$
$avg(|A+FT^{-1}(W)[i,j]|)\simeq av(A)$
Then, it is clear that the denser a graph, the more $A[i,j]$ coefficients are equal to $1$, and decreases the probability of collision (as seen in Equation (C-A)).
C-C Successful Embeddings: a Corollary of Uniqueness
In this section, we show that studying a successful embedding is a corollary of the analytical form of the uniqueness. The embedding of a key $\omega$ in a graph $G$ is successful if the watermarked graph $G_{W}$ is different from the original graph $G$. In other words, $G_{W}$ must be related to its key $\omega$. In fact, the key is binarized after spreading over the entire adjacency matrix. It can happen that the binarization discards a key built with insufficiently large parameters, so that no watermark is inserted into the graph. The study of successful embedding is therefore the estimation of $P(G_{W}=G)$.
Proof
Since $\texttt{Keygen}(0,0)=[]$ (an empty key) and $\texttt{Embed}(G,[])=G$, studying $P(G_{W}=G)$ is a special case of Equation (1) where $G_{W_{1}}=G_{W}$ (the graph watermarked with the key $\omega$) and $G_{W_{2}}=G$ (the graph watermarked with the empty key [][][ ]).
Note that in practice, the owner of the graph knows when the embedding has succeeded or failed: if the watermarked graph is equal to the original graph after an embedding attempt (which is linear to check in $O(|E|)$ with the edit distance for example), it is sufficient for the owner to retry an embedding operation with larger key parameters until the two graphs are truly different. (This will be studied empirically in Section IV-C.)
Appendix D Scalability reduction
In Section IV-B1, we explain that a scability reduction is applied to FFG as it is the case in  [13, 38]. This section provides the details and explanation of such an operation.
When considering graphs, the most significant information is contained in the vertices of the highest degree vertices; thus, analogous to low frequencies, in FFG, the watermark is inserted into the lowest amplitude coefficients in the spectrum among the higher-degree vertices. While the size of images on which the Cox et al. scheme is typically applied is in the order of tens of thousands of pixels, we also choose to take the top $N_{0}=10^{4}$ higher-degree vertices of $G$ to apply FFG onto.
In this way, Embed and Extract are applied only on $N_{0}$ vertices and not on the $N$ vertices of the original graph. Their worst-case time complexities, computed in Section III-D, are now $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. However, to determine the subgraph on which to apply FFG, it is necessary to select the $N_{0}$ vertices of highest degrees (which is calculated in $\mathcal{O}\left(N\log N_{0}\right)$). This cost is added naturally to the Embed function.
The worst-time complexities of FFG, with the dimensionality reduction, are summarized in Table IV.
FFG raw complexities adapted complexities
Keygen $\mathcal{O}(m)$ $\mathcal{O}(m)$
Embed $\mathcal{O}\left(N^{2}\log N\right)$ $\mathcal{O}\left((N+N_{0}^{2})\log N_{0}\right)$
Extract $\mathcal{O}\left(N^{2}\log N\right)$ $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$
TABLE IV: Worst-time complexities of FFG to watermark a graph $G$ with $N$ vertices with a key of length $m$ with and without the dimensionality reduction to $N_{0}$ vertices.
D-A Guideline’s complexities
In Section IV-C, we give practical guidelines for low distortion. We now studied this operation in detail, and its effect on the scalability of FFG.
As mentioned before, we set the key length $m$ of FFG as the key length used in [13]. In fact, $m$ can be set in worst-case complexity $\mathcal{O}(N_{0})$. Once the key length $m$ is set, it is always possible to find a value $\sigma$ to watermark the graph while achieving the distortion goal. Experiments have shown that for any graph $G$ and for any fixed key length $m\in[\![1,N_{0}*(N_{0}-1)/2]\!]$, there exists a value $\sigma_{max}$ to watermark the graph with a strictly positive edit distance. $\sigma_{max}$ can be estimated with at most $\mathcal{O}(\log N_{0})$ operations. An automated way to set $\sigma$ resulting in a small ED guarantee, is a dichotomous search in $[\![1,\sigma_{max}]\!]$. The worst-case time complexity of such setting is testing $\mathcal{O}(\sigma_{max}\log\sigma_{max})$ values for $\sigma$ when each test is of complexity $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. To summarize, the worst-case time complexity for setting $\sigma$ knowing $m$ is $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. This is the same as the worst-case complexity of the embedding function, so the proposed guidelines do not affect the scalability of FFG.
Appendix E Additional experiments
E-A (Goal 5) Testing undetectability w. Graph Neural Networks
In addition to the properties stated in the paper and which follow the properties stated in the related papers, we add a Goal 5. We want to measure whether a graph neural network is able to distinguish a watermarked graph from an original graph. Indeed, if such a network existes, it could be maliciously used by an adversary in addition to the classical edge-flip attack.
We present a novel attack strategy. So far, the approach studied has been to modify the graph in an attempt to remove the watermark. While the two related approaches assume that their scheme is undetectable by construction, we decide to go a step further and propose a new attack that focuses on identifying the watermarked graphs instead. If successful, this implies that the watermark significantly degrades the graph despite a small ED. To achieve this, we use a conventional Graph Neural Network (GNN) that solves a graph classification task. To train a predictor for identifying watermarked graphs, we first need a labeled dataset.
We generated $2,000$ independent Barabási-Albert graphs of $1,000$ vertices with an attachment parameter of $3$. Subsequently, we watermarked half of these graphs by generating $1,000$ independent keys with parameters $m=200$ and $\sigma=100$ (chosen as outlined in Section IV-C to achieve an ED between $1$ and $100$ edges of the graph, i.e. between $0.034$ and $3.4$% edges). Our dataset includes the $1,000$ non-watermarked graphs (of label 0) and the $1,000$ watermarked graphs (of label 1). Notably, we underline that the seed graphs for generating the watermarked graphs were excluded from the dataset, to avoid bias in the classification task.
We adapted the graph classification tutorial from the Python library called PyTorch Geometric [30] with our dataset. The task of the GNN is to distinguish graphs into two classes: those who have been watermarked and those who were not. The degrees of the vertices have been added as their features in this dataset. In other words, the degree sequence of each graph will serve as a basis for attempts in detecting a deviation from the benign sequences of non watermarked graphs. The training is set to $80$% of the dataset, leaving $20$% of graphs as a test set.
Our experiments show that with its default parameters or some classic variants of these (number of hidden channels, learning rate…), the GNN is not able to succeed in the task of discriminating between watermarked graphs and unwatermarked graphs. Indeed the accuracy on train and test datasets oscillates between 0.4775 and 0.5225 depending on the hyperparameters.
We conclude that this attack is ineffective against FFG. This implies that the watermarks inserted in FFG are small enough (ED between $0.034$ and $3.4$% edges) to be undetectable by such a GNN classification attack.
E-B Testing Laplacian eigenvalue’s stability
In our examples, we focus on Edit Distance (ED) due to its generality and widespread use in the state-of-the-art. We also consider degree centrality, as shown in Figure 4. The next section discusses the impact of FFG on the Laplacian eigenvalues of a graph, a metric that is sometimes used in graph studies.
[…]
In conclusion, we suggest that future readers test their own metrics of interest before sharing graphs watermarked with Cox-G. Even if there is no theoretical guarantee on the preserving of this metric, the variation induced for the watermark is sometimes negligible for the desired use.